ABC Corporation just paid an annual dividend of $0.75. Analysts expect this dividend to grow at 10% in 2018, 8% in 2019, 5% in 2020, and then maintain a constant growth rate of 2%. If the expected cost of capital is 7% what is the present value of the stock?
Answer #1

Dividend in year 2018=0.75*(1+10%)=0.825
Dividend in year 2019=0.825*(1+8%)=0.891

Dividend in year 2020=0.891*(1+5%)=0.93555

Terminal value=(Dividend in year 2020)*(1+Constant growth rate)/(Cost of capital - Constant growth rate)
=(0.93555)*(1+2%)/(7% - 2%)=19.08522

Now, we need to discount the values using the cost of capital
Present value=0.825/(1+7%)^1+0.891/(1+7%)^2+0.93555/(1+7%)^3+19.08522/(1+7%)^3
=0.771028037+0.778233907+0.763687479+15.57922457
Present value=$17.89